# Beyond Exact Match: Semantically Reassessing Event Extraction by Large Language Models

### v1

@@ Abstract
Event extraction has gained extensive research attention due to its broad range of applications. However, the current mainstream evaluation method for event extraction relies on token-level exact match, which misjudges numerous semantic-level correct cases. This reliance leads to a significant discrepancy between the evaluated performance of models under exact match criteria and their real performance. To address this problem, we propose RAEE, an automatic evaluation framework that accurately assesses event extraction results at semantic-level instead of token-level. Specifically, RAEE leverages Large Language Models (LLMs) as automatic evaluation agents, incorporating chain-of-thought prompting and an adaptive mechanism to achieve interpretable and adaptive evaluations for precision and recall of triggers and arguments. Extensive experimental results demonstrate that: (1) RAEE achieves a very high correlation with the human average; (2) after reassessing 14 models, including advanced LLMs, on 10 datasets, there is a significant performance gap between exact match and RAEE. The exact match evaluation significantly underestimates the performance of existing event extraction models, particularly underestimating the capabilities of LLMs; (3) fine-grained analysis under RAEE evaluation reveals insightful phenomena worth further exploration. The evaluation toolkit of our proposed RAEE will be publicly released.
Beyond Exact Match: Semantically Reassessing Event Extraction by Large Language Models
Anonymous ACL submission
@@ 1 Introduction
Event extraction (EE) aims to detect event instances as well as all of its participants and attributes by analyzing and identifying mentions of semantically defined entities and relationships within them Grishman (2015); Wei and Wang (2019). It is widely divided into two subtasks, Event Detection (ED) and Event Argument Extraction (EAE). While ED aims to identify all triggers that most explicitly convey the occurrence of an event, EAE extracts arguments with semantic roles corresponding to the trigger. Since EE benefits a broad range of applications Zhang et al. (2020); Han et al. (2021), it has gained extensive research attention in recent years Lin et al. (2020); Lu et al. (2022); Yan et al. (2023). Additionally, as large language models (LLMs) are increasingly used as agents Chen et al. (2024); Liu et al. (2023a), the need for accurate event extraction becomes more important, which helps in effectively tracking and labeling the actions of agents in real-world scenarios Park et al. (2023).
@@
Figure 1: One case of EAE is misjudged by exact match (EM) evaluation but reassessed as correct by our proposed RAEE evaluation. With a given trigger (walk), EAE aims to extract corresponding arguments with semantic roles.
@@
Currently, the most widely-used evaluation method for EE is token-level exact match (EM) [^1] For the sake of brevity, we will refer to exact match evaluation as EM in the following content., which assesses whether the predicted triggers and arguments are identical to the annotated ground-truth in benchmarks Hsu et al. (2022); Lu et al. (2022); Hsu et al. (2023a). However, this EM evaluation often misjudges a significant number of cases that are semantically correct or convey the same meaning as ground-truth but do not match it literally Li et al. (2021). For example, one case of EAE task, which extracts event arguments corresponding to a given trigger, is as shown in Figure 1. The predicted arguments, Path: on the Star Ferry terminal concourse and Direction: east, are not the same as the ground-truth and are evaluated as failure extractions under EM. However, it is obvious to assess them as semantically correct extractions. Also, co-reference instances are misjudged by EM, as shown in the example where Promenade and the past Kowloon - Canton Railway Terminus refer to the same thing. Consequently, models that may appear to perform poorly under EM actually have strong performance, indicating that the capabilities of existing EE models might be significantly underestimated. This issue is particularly pronounced in generative models, especially LLMs Huang et al. (2024), where the generations are often semantically correct.
To address this severe issue, we propose a reliable automatic evaluation framework designed to ReAssess Event Extraction models at semantic-level instead of token-level, named RAEE. Specifically, advanced LLMs are utilized to automatically assess the quality of EE by considering their semantic correctness within their context, such as the source document or sentence. Also, an adaptive prompting mechanism is designed to guide LLMs in generating interpretable and adaptive evaluation results for precision and recall of triggers and arguments, in a chain-of-thought approach Wei et al. (2023).
@@
Meta-evaluation results on 8 models, including advanced LLMs, across 10 datasets demonstrate that RAEE achieves a very high correlation with human judgments and significantly outperforms EM. This indicates that our proposed RAEE is a more reliable EE evaluation framework. Additionally, we conduct a comprehensive reassessment for 14 models on 10 datasets using our proposed RAEE evaluation metrics. Extensive experiments reveal a significant performance gap between EM and RAEE, suggesting that existing baselines, particularly LLMs, have been greatly underestimated. Although LLMs still slightly underperform compared to smaller fine-tuned models under RAEE evaluation, the performance gap has narrowed considerably. Furthermore, the ranking order among these baselines has shifted significantly after reassessing under RAEE on most datasets. For example, on ED task, Tagprime-C Hsu et al. (2023a) and DEGREE Hsu et al. (2022) significantly outperform other baselines that perform better under EM. Finally, a fine-grained analysis of the correct and incorrect cases identified by RAEE uncovers intriguing and insightful phenomena worth future exploration. For example, while most spans are accurately matched, their event and role types are usually incorrect under RAEE evaluation. In summary, the contributions of our study are as follows:
@@
  • RAEE: We propose a reliable evaluation framework for the event extraction task, leveraging advanced LLMs along with chain-of-thought prompting and an adaptive mechanism, demonstrating strong correlations with human judgments.
@@
  • Comprehensive evaluation: We comprehensively reassess the performance of previous works and LLMs for the event extraction task using our proposed RAEE framework. Experimental results reveal that the capabilities of event extraction models are significantly underestimated, especially of generative models and LLMs.
@@
  • Fine-grained analysis: We conduct a fine-grained analysis to uncover the correct and incorrect cases identified by RAEE, highlighting areas worth further exploration.
@@ 2.1 Evaluations on Event Extraction
Current evaluations of EE typically use precision, recall, and f1-score metrics to evaluate the span and the type of extracted instances, and this approach is consistent across ED and EAE. The majority of research relies on EM evaluation Lin et al. (2020); Ma et al. (2022); Hsu et al. (2023b), though some studies employ language toolkits to assist with token-level evaluation Li et al. (2021). Additionally, some work modifies the existing evaluation frameworks by incorporating different averaging methods for calculating f1-score Zheng et al. (2021) or introducing additional metrics from downstream applications Peng et al. (2023a). The consistency and fairness of variable alignment with model outputs are also emphasized in recent studies Peng et al. (2023b); Huang et al. (2024). However, these researches uniformly adopt EM when determining span matches, overlooking instances where the span may not exactly match but is semantically correct. To address this oversight, we propose a semantic-level evaluation framework RAEE and provide a more comprehensive evaluation of model performance.
@@ 2.2 LLMs as Judge Models
There is an emerging trend in the research community to leverage large language models (LLMs) for automatically evaluating the quality of generations, thereby reducing the significant cost of human annotations Lan et al. (2024); Lin et al. (2024); Luo et al. (2023); Wang et al. (2023); Liu et al. (2023b); Fu et al. (2023). Extensive studies Fu et al. (2023); Wang et al. (2023) have shown that advanced LLMs, such as GPT-4, achieve very close correlations with human judgments across various generative NLP tasks, including dialogue generation Fu et al. (2023), text summarization Liu et al. (2023b) and structured generation Wang et al. (2023).
In this paper, to address the limitations of EM, we utilize advanced LLMs to mimic human annotators in automatically evaluating the quality of model-generated trigger words and arguments for EE task. Advanced LLMs achieve performance that is closely aligned with human judgments while incurring significantly lower costs Lan et al. (2020), making our proposed RAEE a scalable and reliable automatic evaluation method for EE task.
3 Automatic Reassessing Framework RAEE
To address the misjudgments caused by EM, we propose an automatic evaluation framework, RAEE, to reassess existing event extraction methods more accurately. In this section, we will introduce: (1) how RAEE differs from EM in Section 3.1; (2) how RAEE automatically achieves interpretable and adaptive evaluations in Section 3.2; (3) the reliability of RAEE in Section 3.3.
@@ 3.1 Semantic-level F1-score
Figure 2 shows the difference between RAEE and EM. In EM, True Positive (TP) cases are identical in the calculation of precision and recall. While in RAEE, we design the semantic-level F1-score by calculating precision and recall separately without counting TP as follows:
$\bf p=\frac{|P_{c}|}{|P|},\bf\quad r=\frac{|G_{r}|}{|G|},\quad\bf f1_{\rm RAEE}=\frac{2*p*r}{p+r}$
where $\bf P$ and $\bf G$ are predicted results and golden annotations. $\bf P_{c}$ and $\bf G_{r}$ are correct predicted results and recalled golden annotations, respectively.
Figure 2: Comparison of EM and RAEE evaluation using EAE as an example (consistent to ED). RAEE utilizes a LLM to assess event extraction results from both precision and recall perspectives. The source of the arrows indicates the basis or reason for judgment, with solid arrows representing the precision evaluation process and dashed arrows representing the recall evaluation process.
Since human evaluation is the most reliable golden standard Wang et al. (2023); Fu et al. (2023), the most accurate results for $\bf P_{c}$ and $\bf G_{r}$ should ideally be obtained through human annotation. However, due to the high cost, lengthy duration, and lack of reproducibility associated with human evaluation Lan et al. (2024, 2020), the advanced LLM, GPT-4o, is utilized as a reliable proxy for human evaluation in RAEE. Specifically, LLMs are equipped with an adaptive mechanism and chain-of-thought prompting strategies to evaluate the quality of event extraction results by considering the semantic correctness within context.
3.2 Adaptive Evaluation by LLMs
Human annotators might have slightly different subjective perspectives when evaluating, which could lead to differences in assessment Lan et al. (2024); Kim et al. (2023). Therefore, for both ED and EAE tasks, we manually identify several key criteria of evaluation bias, like evaluating pronouns in co-reference as correct extractions, and prompt LLMs to follow them to judge the quality of event extraction results step by step. More details about our human-annotated key criteria of evaluation bias and reassessment prompts can be found in Appendix C and Appendix D.1, respectively.
It should be noted that these key criteria are optional and adaptive, contributing to the adaptive mechanism in prompts. This allows researchers to tailor the evaluation criteria of RAEE to their specific needs in different scenarios. For example, some datasets provide separate annotations for co-references to assist in the annotation of event extraction Li et al. (2021) but some do not Ebner et al. (2020); Satyapanich et al. (2020). To achieve fair comparison in experiments, the evaluation criteria remain the same for all the evaluated models in our work.
Task - Agree. $\uparrow$ Spearman $\uparrow$
ED-p Human-Average 89.08 74.18
GPT4o-Human 76.89 49.09
      - w/o. Adaptive 78.99 47.25
EM-Human 26.05 ——
ED-r Human-Average 96.87 92.64
GPT4o-Human 87.17 74.72
      - w/o. Adaptive 84.19 69.75
EM-Human 71.05 ——
EAE-p Human-Average 93.61 85.47
GPT4o-Human 88.13 72.55
      - w/o. Adaptive 77.28 53.16
EM-Human 63.49 ——
EAE-r Human-Average 96.96 94.00
GPT4o-Human 86.10 75.27
      - w/o. Adaptive 86.42 76.08
EM-Human 72.55 ——
Table 1: Agreement percent and Spearman correlation with human average. -p and -r represent the evaluation of precision and recall, respectively. All Spearman correlation scores are associated with the $p$-value $<$ 0.05, indicating the results are statistically significant. Experimental results denote that LLM achieves a moderate and strong correlation with human judgments. Since EM fails to achieve basic correlation scores with human judgments, their Spearman correlation scores cannot be computed correctly.
3.3 RAEE Correlation with Human Evaluation
To prove the effectiveness of our proposed RAEE evaluation framework and adaptive evaluation, we conduct the meta-evaluation to analyze the correlation between RAEE and human judgments. Specifically, we invite 3 professional event extraction annotators to evaluate 200 sampled items of data for both ED and EAE tasks on precision and recall metrics. These annotators are given the same prompts as those given to LLM. With reassessing results from 3 humans and a LLM (GPT-4o), we calculate two widely-used inter-annotator agreement (IAA) metrics among humans and LLM: percent agreement Zheng et al. (2023); Liu et al. (2021) and Spearman correlation Lan et al. (2020, 2024); Fu et al. (2023). The experimental results are shown in Table 1.
3.3.1 Correlation on Recall Task
While LLM correlation with human average shows a noticeable drop (average 18.33 Spearman and 10.28% agreement percent), they still achieve a very strong correlation with human judgments on Agreement percent and Spearman correlation, i.e., over 86% agreement percent, which significantly outperforms EM with over 14.84% agreement percent. This also suggests that while the LLM can simulate human judgments and be regarded as an automatic evaluation agent, there is still room for further refinement.
3.3.2 Correlation on Precision Task
LLMs achieve a strong correlation with human judgments, over 76% agreement percent and 0.49 Spearman correlation of ED-p Lan et al. (2020); Fu et al. (2023). Furthermore, the average correlation scores between LLMs and human average for ED-p are significantly lower than that in the other tasks, indicating that semantic evaluation in this task is more challenging. Our investigation reveals two main reasons for this phenomenon: (1) Within the same domain, different datasets have personalized definitions for event types with similar names; (2) As for datasets across different domains, there are discrepancies in how events are defined as being triggered. When sampling data from multiple datasets, these subtle definitional differences can lead to confusion among evaluators when determining whether trigger predictions are accurate, leading to lower consistency in ED-p. In contrast, evaluating recall is simpler since it only requires determining whether gold annotations are retrieved, without judging the accuracy of the event type. Also, evaluating event arguments is more straightforward, as the role types for event arguments are clearly defined across datasets. This explains the relatively higher IAA observed in the three tasks other than ED-p.

### v2

@@ Abstract
Event extraction has gained extensive research attention due to its broad range of applications. However, the current mainstream evaluation method for event extraction relies on token-level exact match, which misjudges numerous semantic-level correct cases. This reliance leads to a significant discrepancy between the evaluated performance of models under exact match criteria and their real performance. To address this problem, we propose a reliable and semantic evaluation framework for event extraction, named RAEE, which accurately assesses extraction results at semantic-level instead of token-level. Specifically, RAEE leverages large language models (LLMs) as evaluation agents, incorporating an adaptive mechanism to achieve adaptive evaluations for precision and recall of triggers and arguments. Extensive experiments demonstrate that: (1) RAEE achieves a very strong correlation with human judgments; (2) after reassessing 14 models, including advanced LLMs, on 10 datasets, there is a significant performance gap between exact match and RAEE. The exact match evaluation significantly underestimates the performance of existing event extraction models, and in particular underestimates the capabilities of LLMs; (3) fine-grained analysis under RAEE evaluation reveals insightful phenomena worth further exploration. The evaluation toolkit of our proposed RAEE is publicly released [^1] https://github.com/Lyfralston/RAEE.
Beyond Exact Match: Semantically Reassessing Event Extraction
by Large Language Models
Yi-Fan Lu1, Xian-Ling Mao1, Tian Lan1, Heyan Huang1, Chen Xu1, Xiaoyan Gao2 1Beijing Institute of Technology, 2Beijing University of Technology yifanlu@bit.edu.cn, maoxl@bit.edu.cn, lantiangmftby@gmail.com, chenxu05037@gmail.com, hhy63@bit.edu.cn, gaoxiaoyan@bjut.edu.cn
@@ 1 Introduction
Figure 1: One case of EAE is misjudged by exact match (EM) evaluation but reassessed as correct by our proposed RAEE evaluation. With a given trigger (walk), EAE aims to extract its arguments with semantic roles.
Event extraction (EE) aims to detect event instances as well as all of their participants and attributes by analyzing and identifying mentions of semantically defined entities and relationships within them Grishman (2015); Wei and Wang (2019). It is typically divided into two subtasks, Event Detection (ED), which focuses on identifying triggers that indicate event occurrence, and Event Argument Extraction (EAE), which extracts arguments with semantic roles related to the trigger. Since EE benefits a broad range of applications Zhang et al. (2020); Han et al. (2021), it has gained extensive research attention Lu et al. (2022); Yan et al. (2023). With the increasing use of large language models (LLMs) as agents Chen et al. (2024b); Liu et al. (2023a), the requirement for accurate event extraction becomes more important, which helps in effectively tracking and labeling the actions of agents in real-world scenarios Park et al. (2023).
@@
Currently, the most widely-used evaluation method for EE is token-level exact match (EM) [^2] For the sake of brevity, we will refer to exact match evaluation as EM in the following content., which assesses whether the predicted triggers and arguments are identical to the annotated ground-truth in benchmarks Hsu et al. (2022); Lu et al. (2022); Hsu et al. (2023a). However, EM often misjudges a significant number of cases that are semantically correct or convey the same meaning as ground-truth but cannot match it literally Li et al. (2021). For example, one case of EAE task, which extracts event arguments corresponding to a given trigger, is illustrated in Figure 1. The predicted arguments, Path: on the Star Ferry terminal concourse and Direction: east, are not the same as ground-truth and are therefore evaluated as failed extractions under EM. However, considering the context, both predictions can be regarded as semantically correct extractions. Also, co-reference instances are misjudged by EM, as shown in the example where Promenade and the past Kowloon - Canton Railway Terminus refer to the same place. Consequently, models that may appear to perform poorly under EM actually have strong performance, indicating that the capabilities of existing EE models might be greatly underestimated. This issue is particularly prominent in generative models, especially LLMs Huang et al. (2024), where the generations often cannot exactly match ground-truth.
To address this severe issue, we propose a reliable and semantic evaluation framework designed to ReAssess Event Extraction models at semantic-level instead of token-level, named RAEE. Specifically, RAEE utilizes advanced LLMs as evaluation agents to achieve automatic assessment. By considering the semantics of extractions in the context, the LLM performs binary judgments on the extractions from the perspectives of precision and recall, determining whether they are corrected and whether they are successfully recalled, respectively. To ensure the reliability of the evaluation, we introduce an adaptive prompting mechanism to enhance LLMs’ evaluation consistency with human evaluation.
@@
Experiments begin with a meta-evaluation designed to validate evaluation reliability of RAEE. The results demonstrate that (1) RAEE achieves a strong correlation with human judgments, (2) RAEE significantly outperforms EM and other soft matching evaluation methods, and (3) the proposed adaptive mechanism enhances the reliability of RAEE. Then, we apply RAEE to conduct a comprehensive reassessment experiment for 14 EE models on 10 widely-used datasets. Extensive experiments reveal that (1) evaluated models exhibit a significant performance gap under EM and RAEE, indicating that existing EE models, particularly generative models, are greatly underestimated, and (2) the ranking order among these models shifts after reassessing under RAEE on most datasets, further highlighting the importance of RAEE. To further explore the reasons behind misjudgments evaluated by EM and failure modes evaluated by RAEE, a fine-grained analysis experiment uncovers insightful phenomena worth future exploration. For example, while most spans are accurately matched, their types are still incorrect under RAEE evaluation. Finally, we reassess LLMs’ event extraction ability and find that LLMs are extremely underestimated by EM, further emphasizing the urgency of incorporating semantic evaluation in the era of LLMs. In summary, our contributions are as follows:
@@
  • RAEE: We propose a reliable and semantic evaluation framework to assess event extraction at semantic-level by leveraging advanced LLMs as evaluation agents with an adaptive prompting mechanism. RAEE shows strong correlations with human judgments.
@@
  • Comprehensive evaluation: We apply our proposed RAEE to comprehensively reassess the performance of previous works and LLMs for event extraction. Experimental results reveal that the capabilities of event extraction models are significantly underestimated, especially of generative models and LLMs.
@@
  • Fine-grained analysis: We conduct a fine-grained analysis to uncover the reasons behind misjudgments evaluated by EM and failure modes evaluated by RAEE, highlighting areas worth further exploration.
Figure 2: Evaluation process of RAEE, using the precision of EAE as an example (consistent with ED).
@@ 2.1 Evaluations on Event Extraction
Current evaluations of EE typically use precision, recall, and f1-score metrics to evaluate the span and the type of extracted instances, and this approach is consistent across ED and EAE. The majority of research relies on EM evaluation Lin et al. (2020); Ma et al. (2022); Hsu et al. (2023b), and some studies employ language toolkits to apply head noun evaluation Li et al. (2021). Additionally, some works modify the existing evaluation frameworks by incorporating different averaging methods for calculating f1-score Zheng et al. (2021) or introducing additional metrics from downstream applications Peng et al. (2023a). The consistency and fairness of variable alignment with model outputs are also emphasized in recent studies Peng et al. (2023b); Huang et al. (2024). However, these works uniformly adopt EM when determining span matches, and overlook instances where the semantically correct span cannot exactly match ground-truth. To address this oversight, we propose a semantic-level evaluation framework and provide a more reliable evaluation of model performances.
@@ 2.2 LLMs as Judge Models
There is an emerging trend in the research community to leverage LLMs for automatically evaluating the quality of generations, thereby reducing the significant cost of human annotations Lan et al. (2024); Lin et al. (2024); Luo et al. (2023); Fu et al. (2023); Wang et al. (2023). Extensive studies Fu et al. (2023); Wang et al. (2023) have shown that advanced LLMs, such as GPT-4, achieve very close correlations with human judgments across various generative NLP tasks, including dialogue generation Fu et al. (2023), text summarization Liu et al. (2023b) and structured generation Wang et al. (2023).
In this paper, we utilize advanced LLMs to emulate human annotators in automatically evaluating EE. Advanced LLMs closely align with human judgments at a fraction of the cost Lan et al. (2020), making our RAEE a scalable and reliable automatic evaluation method for EE.
3 Semantic Reassessing Framework
To address the misjudgments caused by EM, we propose a semantic evaluation framework, named RAEE. In this section, we will introduce: (1) how semantic-level F1-score is calculated in RAEE in Section 3.1; (2) the adaptive mechanism of RAEE in Section 3.2; (3) the evaluation reliability of RAEE in Section 3.3.
@@ 3.1 Semantic-level F1-score
Assume there are a piece of context $\bf C$, predicted results from an EE model ${\bf P}={\rm Model}(\bf C)$, and gold annotations from human ${\bf G}={\rm Human}(\bf C)$. RAEE prompts an evaluation agent LLM to consider the context during evaluation, assessing predictions and annotations separately when calculating precision and recall. Specifically, RAEE requires the LLM to identify a correct prediction set $\bf P_{c}$ and a recalled annotation set $\bf G_{r}$:
${\bf P_{c}}={\rm LLM}({\bf P|C,G}),\quad{\bf G_{r}}={\rm LLM}({\bf G|C,P}).$
Then, semantic-level f1-score can be computed as follows:
$\bf p=\frac{|P_{c}|}{|P|},\bf\quad r=\frac{|G_{r}|}{|G|},\quad\bf f1_{\rm RAEE}=\frac{2*p*r}{p+r}.$
Figure 2 presents an example to evaluate predictions ${\bf P}$ in EAE. The LLM is required to first analyze the context and then evaluate each prediction with a reason to get correct predictions $\bf P_{c}$.
Eval. Methods ED precision ED recall EAE precision EAE recall
Per. Sp. Per. Sp. Per. Sp. Per. Sp.
Human Average 89.08 74.19 96.87 92.65 93.61 85.48 96.96 94.00
Deepseek-r1 77.451.65 53.192.57 86.850.22 72.310.54 79.720.42 55.170.78 93.490.53 87.021.08
      - w/o. Adaptive 72.600.93 48.211.16 85.240.20 71.130.33 76.980.51 54.241.00 91.750.41 83.470.82
GPT-4o-240513 78.290.50 50.492.31 86.850.80 71.701.74 84.630.69 64.932.27 83.710.55 68.810.54
      - w/o. Adaptive 75.301.41 38.333.80 83.721.23 69.331.92 80.561.32 57.732.63 82.562.61 66.885.00
Gemini-1.5-pro 72.830.89 37.932.26 51.490.77 34.600.73 81.350.61 55.031.48 77.160.22 61.810.44
      - w/o. Adaptive 71.941.06 28.143.12 48.511.77 31.192.21 79.570.89 49.622.59 68.551.42 49.742.37
Deepseek-v3 66.391.59 28.613.38 71.472.51 54.033.13 77.592.39 52.784.41 76.330.54 57.061.55
      - w/o. Adaptive 66.711.49 31.071.91 61.872.95 44.043.37 72.830.28 48.771.71 72.101.93 49.872.52
Claude-3.5-sonnet 63.631.09 33.982.96 71.620.92 51.771.30 75.610.57 48.651.31 71.881.01 50.751.29
      - w/o. Adaptive 61.480.52 34.831.47 64.323.04 43.494.33 68.910.19 44.411.03 68.621.43 43.953.09
Similarity 50.70 12.62 84.82 71.40 63.81 16.76 75.79 54.44
Head Noun 31.93 11.80 75.74 29.97 40.98 20.07 64.39 35.98
Exact Match 28.57 —— 72.46 —— 31.28 —— 54.51 ——
Table 1: Percent agreement (%) and Spearman correlation ($\times$100) of different evaluation methods with human judgments on four event extraction subtasks. All LLM experiments are repeated three times. All Spearman correlation scores are associated with the $p$-value $<$ 0.05. For each metric in each subtask, we highlight the highest value using a dark shade, and the second-highest value using a light shade.
3.2 Adaptive Mechanism
Human annotators may have slightly different subjective perspectives when performing semantic evaluation, which could lead to inconsistencies in assessment Lan et al. (2024); Kim et al. (2023). Besides, various datasets often have customized annotation rules, resulting in differing evaluation criteria. For example, some datasets provide separate annotations for co-references to assist in the annotation of event extraction Li et al. (2021) but some do not Ebner et al. (2020); Satyapanich et al. (2020). To mitigate inconsistent evaluation and achieve adjustable evaluation in RAEE, we manually identify several key criteria of evaluation bias, and prompt the LLM to follow them to assess extractions step by step, as illustrated in Figure 2. When an evaluated content meets a specific evaluation criteria, the evaluation agent will explicitly output the corresponding criteria it follows.
These key criteria are adaptive, which allows researchers to tailor the evaluation criteria of RAEE to their specific needs in different scenarios. In the following evaluation experiments, these key criteria remain consistent to achieve consistent evaluation. The default setting of existing criteria can be found in Appendix B.
3.3 Reliability of RAEE
To validate the reliability of RAEE and the effectiveness of its adaptive mechanism, we conduct a meta-evaluation to analyze the correlation between RAEE and human judgments. Specifically, we sample 800 pieces of data in ED and EAE tasks from mainstream event datasets (presented in the main experiment) across various domains, excluding extractions easily verified as correct or recalled by EM. We apply different LLMs as the evaluation agent to investigate the impact of agent selection on RAEE. The full prompts are provided in Appendix C.1. For human judgments, 3 professional event extraction annotators evaluate the same data using the same prompts as the LLMs. We also apply EM, head noun evaluation Li et al. (2021) and text similarity evaluation with a well-selected threshold $t=0.5$ Chen et al. (2023) for comparison. Two widely-used inter-annotator agreement (IAA) metrics are selected: Percent agreement Zheng et al. (2023); Liu et al. (2021) and Spearman correlation Lan et al. (2024); Fu et al. (2023). The results are presented in Table 1.
3.3.1 Overall Correlation
When using deepseek-r1 as the evaluation agent of RAEE, although there is a noticeable gap between RAEE and human average evaluation (average 0.197 Spearman correlation and 9.75% Percent agreement), RAEE still shows strong consistency with human judgments [^3] These metrics reflect a strong correlation, as validated in past works Lan et al. (2020); Fu et al. (2023)., achieving over 77% Percent agreement and 0.531 Spearman correlation. Compared to traditional evaluation methods, RAEE significantly outperforms exact match and head noun evaluations, improving Spearman correlation by over 0.351 across all tasks. Notably, the similarity-based evaluation shows little difference from RAEE on the ED recall task. This is because golden triggers to be evaluated in ED recall task are correct and typically single-word mentions. When the evaluation target shifts to model predictions that might be incorrect, as in the precision task, or involves multi-word mentions, as in the EAE task, the performance of similarity-based evaluation significantly declines.
Furthermore, for all evaluation methods, the correlation with human judgments in precision tasks are significantly lower than those in recall tasks. This discrepancy arises from the inherent difficulty in evaluating precision compared to recall. Specifically, in the precision evaluation task, the evaluation focuses on assessing the correctness of model predictions. However, since the evaluated targets in the recall evaluation task are assumed to be correct, the evaluation primarily involves identifying potential correspondences. This phenomenon is also observed among human evaluators, as the correlation among humans is lower in the precision task compared to the recall task.